Three consecutive even integers are such that the square of the third is 76 more than the square of the second. Find the three integers.
Answer by checkley75(3666)

(x+4)^2=(x+2)^2+76
X^2+8x+16=x^2+4x+4+76
8x-4x=80-16
4x=64
x=64/4
x=16 answer for the first number.
16+2=18 for the middle number .
16+4=20 for the highest number.
proof
20^2=18^2+76
400=324+76
400=400